# Non-thermal X-ray Emission from Merging Massive Black Hole Binaries

### v1

@@ 4 Conclusions and Discussion
In this work, we have investigated whether non-thermal X-ray emission powered by magnetic reconnection could obscure or modify the recently identified disappearing thermal X-ray signature associated with merging massive black hole binaries embedded in circumbinary disks. Motivated by hydrodynamical simulations showing a sharp decline in thermal X-ray emission from tidally truncated minidiscs near merger (Krauth et al., 2023), we constructed semi-analytic models for both the thermal and non-thermal components and compared their behaviour across the MBHB mass range relevant for LISA sources.
Our primary result is that, for physically motivated values of the magnetic field strength and accretion rate, the non-thermal synchrotron X-ray luminosity remains several orders of magnitude below the thermal X-ray emission throughout the inspiral. Even when these parameters are pushed to optimistic or extreme values, the non-thermal component remains suppressed by at least $\sim$2 orders of magnitude relative to the thermal emission. This indicates that the disappearance of the thermal X-ray signal identified by Krauth et al. (2023) is unlikely to be masked by reconnection-powered synchrotron radiation in realistic astrophysical systems.
@@
While our analysis is designed to capture the order-of-magnitude evolution of the band-integrated thermal X-ray luminosity, it does not aim to resolve changes to the instantaneous spectral shape during the late inspiral and merger. In particular, magnetic reconnection can accelerate particles into hard power-law distributions, producing non-thermal synchrotron emission that extends into the hard X-ray band, with additional modification at the highest energies possible through Compton upscattering. Because power-law spectra decline more gradually with energy than thermal emission, such non-thermal components could in principle dominate the highest-energy X-ray bands even if they contribute only a small fraction of the total X-ray output. The relative importance of such components depends on the detailed spectral shapes of the thermal and non-thermal emission, and on how they evolve during the final stages of inspiral, which we defer to future work.
@@
Taken together, these results strengthen the robustness of the disappearing thermal X-ray signature as an electromagnetic counterpart to MBHB mergers. The lack of a long-lived non-thermal contaminant implies that the X-ray-dark phase should provide a clean temporal marker associated with the final stages of inspiral and merger. Such a signature could play an important role in identifying host galaxies within the LISA localization volume and in coordinating multi-messenger follow-up observations.
@@
Nevertheless, within the range of parameters explored here, our results indicate that magnetic reconnection is unlikely to erase or substantially alter the disappearing thermal X-ray signature. Instead, both thermal and non-thermal emission are expected to fade rapidly around merger, reinforcing the picture of a brief but distinct X-ray-dark phase preceding the post-merger re-establishment of accretion. This behaviour underscores X-ray observations as a powerful complement to gravitational-wave detections in probing the final stages of massive black hole binary evolution.

### v2

@@ 4 Conclusions and Discussion
In this work, we have investigated whether non-thermal X-ray emission powered by magnetic reconnection could obscure or modify the recently identified disappearing thermal X-ray signature associated with merging massive black hole binaries embedded in circumbinary disks. Motivated by hydrodynamical simulations showing a sharp decline in thermal X-ray emission from tidally truncated minidiscs near merger (Krauth et al., 2023), we constructed semi-analytic models for both the thermal and reconnection-powered non-thermal components. We compared their mass scalings, spectral shapes, and time-dependent evolution across soft and hard X-ray bands over the LISA-relevant MBHB mass range, and incorporated a magnetospheric balding prescription to model the post-merger decay of the non-thermal emission.
Our primary result is that, for physically motivated magnetic field strengths and accretion rates, the non-thermal synchrotron X-ray luminosity remains subdominant to the thermal X-ray emission throughout the inspiral. In the soft X-ray band—where the thermal emission is concentrated and where most X-ray instruments operate—the synchrotron component remains below the thermal level even under optimistic assumptions. When integrated over the full X-ray range, the non-thermal luminosity can approach or exceed the thermal component only at the highest black hole masses and only when compared to a non-spinning ($a=0$) thermal model. For rapidly spinning black holes ($a=0.99$), which are expected to produce the strongest magnetospheres and thus the largest synchrotron output in our framework, the thermal emission remains dominant across the full mass range. This indicates that the disappearance of the thermal X-ray signal identified by Krauth et al. (2023) is unlikely to be masked by reconnection-powered synchrotron radiation in realistic astrophysical systems.
@@
In addition to band-integrated comparisons, our spectral modelling shows that the relative dominance of thermal and non-thermal emission is energy-dependent and evolves across inspiral. At early times the soft X-ray band is clearly thermal-dominated, while toward merger the thermal component collapses and the non-thermal synchrotron emission becomes increasingly important at higher photon energies. However, this transition occurs only after the thermal luminosity has already declined substantially, and the non-thermal component remains insufficient to erase the characteristic soft X-ray disappearance identified by Krauth et al. (2023).
The spectral reconfiguration towards harder photon energies near merger therefore points to a complementary high-energy signature rather than a competing one. A transient enhancement in the hard X-ray band could trace the increasing role of magnetospheric reconnection during the final stages of inspiral. If detectable in nearby systems with favourable combinations of mass, magnetic field strength, and accretion state, such hard-band emission would not obscure the thermal-drop scenario, but instead mark the system’s transition toward a magnetically dominated state, providing an additional observational handle on the approach to merger.
@@
Taken together, these results reinforce the disappearing thermal X-ray signature as a robust electromagnetic counterpart to MBHB mergers. Because the reconnection-powered component does not persist at a level capable of masking the thermal drop, the merger-associated X-ray–dark phase remains a distinct and temporally localized feature suitable for identifying host galaxies within the LISA localization volume and performing multi-messenger follow-up observations. Any accompanying hard X-ray enhancement, if detected, would not contradict this picture, but instead provide complementary information about the onset of magnetospheric dominance marking the final stages of inspiral. In combination, the soft-band disappearance and possible hard-band emergence could therefore offer a multi-band electromagnetic framework for temporally associating LISA detections with their astrophysical environments and advancing broader multi-messenger science goals.
@@
Nevertheless, within the range of parameters explored here, our results indicate that magnetic reconnection is unlikely to erase or substantially alter the disappearing thermal X-ray signature. The spectral evolution further demonstrates that while non-thermal emission can become prominent at higher photon energies near merger—particularly in extreme magnetic-field configurations—the soft X-ray disappearance associated with minidisc depletion remains robust. Although the non-thermal component may briefly rise as magnetic stresses intensify, it subsequently decays on short balding timescales once the external magnetic flux supply diminishes. As a result, the system is expected to enter a transient X-ray–dark phase before accretion and magnetic flux are re-established. This behaviour underscores X-ray observations as a powerful complement to gravitational-wave detections in probing the final stages of massive black hole binary evolution.